OpenStudy (anonymous):

Solve: −2(x − 5) = −4

OpenStudy (mathstudent55):

First distribute the -2 on the left side: -2(x - 5) = -4 -2x + 10 = -4

OpenStudy (mathstudent55):

Now subtract 10 from both sides. Then divide both sides by -2.

OpenStudy (anonymous):

that doesnt explain anything

OpenStudy (mathstudent55):

Subtracting 10 from both sides: -2x + 10 = -4 -10 -10 --------------- -2x = -14 Now divide both sides by -2: -2x -14 ---- = ----- -2 -2 x = 7

OpenStudy (mathstudent55):

Does that help?

OpenStudy (anonymous):

thnx

OpenStudy (mathstudent55):

ur wlcm